OpenStudy (anonymous):

Am I right? :c Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. 2x – y = –7

OpenStudy (uri):

y=mx+b use this formula.

OpenStudy (anonymous):

..I already did all of the calculations, I'd just like to know if I'm right unless you want my calculations and such.

OpenStudy (anonymous):

yeah thats correct

OpenStudy (anonymous):

y-y1=m(x-x1) point slope formula

OpenStudy (anonymous):

thank you c:

OpenStudy (anonymous):

you welcome